$2,000 is deposited today into a bank account. The account earns 4.3% per annum compounded quarterly for the first 4 years, then 6.3% per annum compounded monthly thereafter. Assuming no further deposits or withdrawals are made,

(a) Calculate the account balance six months from today.

(b) Calculate the account balance 4 years from today.

(c) Calculate the account balance 4.25 years from today.

(d) Calculate the account balance 13 years from today.
Answer #1
(a) Account balance six months from today $ 2,043.23
Working:
A = P*(1+i)^n Where,
= 2000*(1+0.01075)^2 P = $ 2,000.00
= $ 2,043.23 i = 4.3%/4 = 0.01075
n = 2
(b) Account balance 4 years from today $ 2,373.18
Working:
A = P*(1+i)^n Where,
= 2000*(1+0.01075)^16 P = $ 2,000.00
= $ 2,373.18 i = 4.3%/4 = 0.01075
n = 16
(c) Account balance 4.25 years from today $ 2,410.75
Working:
A = P*(1+i)^n Where,
= 2373.18*(1+0.00525)^3 P = $ 2,373.18
= $ 2,410.75 i = 6.3%/12 = 0.00525
n = 3
(d) Account balance 13 years from today $ 4,243.78
Working:
A = P*(1+i)^n Where,
= 2410.75*(1+0.00525)^108 P = $ 2,410.75
= $ 4,243.78 i = 6.3%/12 = 0.00525
n = 108